calculate the number of tonnes of CO2 equivalent of 80 kg of N2O


1
Expert's answer
2021-09-06T02:38:01-0400

M(N2O)=44.013 g/mol

n(N2O) "= \frac{80 \times 10^6 \;g}{44.013 \;g/mol}= 1.8176 \times 10^6 \;mol"

M(CO2) = 44.01 g/mol

m(CO2) "= 44.01 \times 1.8176 \times 10^6"

"= 79.99 \times 10^6 \; g = 79.99 \;kg = 0.07999 \;tonnes"

Answer: 0.079 tonnes